Find the hypotenuse of a right triangle with legs of 15 inches and 20 inches.
1
Expert's answer
2013-04-04T04:35:27-0400
The Pythagorean theorem : the sum of the squared legs equals the square of the hypotenuse.
A^2+B^2=C^2
Substituting given values we obtain:15*15+20*20=C^2
C^2=625
Thus, the answer is C=25